OpenStudy (anonymous):

which property is illustrated by the equation -8+0=-8

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge a+0=0+a=a$$

OpenStudy (skullpatrol):

When 0 is added to any number the sum is identical to that number.

OpenStudy (skullpatrol):

$$\Huge\color{blue}{identical }$$

OpenStudy (skullpatrol):

where 0 is the identity element for addition